# Finite Element Simulation of Modified Poisson–Nernst–Planck/Navier–Stokes Model for Compressible Electrolytes under Mechanical Equilibrium

### v3

@@ Abstract
journal: Applied Numerical Mathematics
@@ 1.1  Scope
The ability of an electrolyte to compress or expand is not merely a mathematical nuance but a critical physical factor that governs ion transport and storage in advanced electrochemical devices. Our work, which introduces a total concentration as a spatially variable field, is directly motivated by these applications. Compressibility effects are particularly important in porous electrodes used in supercapacitors and batteries, where nanoconfinement leads to ion saturation and strong local density deviations from the bulk [61], a phenomenon that incompressible models cannot capture. Similarly, ionic liquids and highly concentrated electrolytes represent a class of materials where steric repulsions and ion correlations dominate their electrochemical response. The seminal works of [62] and [63] demonstrated that the finite size of ions imposes a limit on the maximum local density, leading to a compressible medium whose structure under an electric field gives rise to characteristic, non-monotonic capacitance curves. Moreover, in soft matter systems such as polymer electrolytes and biological media [64], the electrolyte resides within a flexible matrix that can swell or contract, coupling mechanical deformation with ion transport, rendering the incompressibility assumption physically invalid in such applications.
@@ 5.3  Symmetric 1:1 electrolyte: Transition to incompressible behavior with increasing Bulk modulus
The physical implications of this density variation are profound and directly impact the electrolyte’s ability to store charge. As discussed in Section 5.7, the enhanced ionic accommodation in compressible systems at high potentials, evident in Fig. 7, is the very mechanism that boosts the double-layer capacitance, $C_{dl}$. Conversely, as $K$ increases and the system becomes incompressible, the number density is constrained near $n^{0}$, reducing the space charge distribution and consequently lowering the overall $C_{dl}$. Therefore, compressible electrolytes are critical in applications [61, 62, 63] where density changes actively influence ion transport and energy storage.
@@ 5.4  Ionic distributions in annular battery designs at equilibrium
Figure 8: Simulation of electrolyte flow on full domain (left) and cross-section along $y=0$, $1\leq x\leq 2$ (right) in an annular battery design in Example 5.4, with an applied potential of 1.0-1.0 unit on the inner boundary and $1.0$ unit on the outer boundary. Parameter settings are provided in Table 3. Boundary conditions are described in Section 3.2.
@@ 5.7  $K$-dependent double layer capacitance curves
The results show that the total electrostatic charge stored in the system increases monotonically as $\varphi^{\mathrm{Left}}$ is varied from 0 to 80 for all values of the compressibility parameter $K$. As $K$ is increased, the total charge decreases and approaches the value corresponding to the incompressible mixture, in agreement with theoretical expectations. This behavior confirms that increasing compressibility reduces the overall charge accumulation in the double layer. Also, near the electrodes, one ionic species becomes strongly depleted due to the applied potential, resulting in a negligible contribution of that species to the total charge. Consequently, the double layer behavior in this region is primarily governed by the counter-ions that accumulate close to the electrode surface.
The differential double layer capacitance, $C_{dl}$, exhibits a similar dependence on the compressibility parameter $K$, but shows a distinct behavior with respect to the applied potential. The observed behavior connects directly to the physical principles established in studies of porous electrodes [61] and ionic liquids [62, 63], where density variations fundamentally govern electrochemical response. It initially increases with $\varphi^{\mathrm{Left}}$, reaches a slight local maximum, then decreases marginally before rising again at higher potentials. This subtle non-monotonic behavior reflects a shift in charge accumulation due to enhanced volume changes in the electrolyte, which allow additional ionic accommodation and thus boost the capacitance at high fields. As $K$ increases, the overall magnitude of $C_{dl}$ decreases and approaches the incompressible limit, indicating that more compressible systems have a reduced ability to store charge within the double layer. Physically, these trends arise because, as discussed in Section 5.3, increasing compressibility reduces the number density toward its reference value, thereby directly affecting the space charge distribution, which also decreases correspondingly. Hence, the results demonstrate that compressible electrolytes are essential wherever density changes influence ion transport and charge storage capabilities.
@@ Acknowledgments
Ankur acknowledges the funding through the Advanced Research Opportunities Program (AROP), Germany and the University Grant Commission, India, ID. No. JUNE18-416131. S.S. is funded by German Research Foundation (DFG) under Research Unit FOR5409: ”Structure-Preserving Numerical Methods for Bulk- and Interface-Coupling of Heterogeneous Models (SNuBIC)”. The authors also sincerely thank Prof. Manuel Torrilhon (Head of the ACoM Lab, RWTH Aachen University, Germany) for his valuable guidance and intellectual input throughout the entire AROP project carried out by Ankur at the ACoM Lab.

### v4

@@ Abstract
journal: Applied Numerical Mathematicsjournal: Applied Numerical Mathematics11footnotetext: Accepted for publication in Applied Numerical Mathematics.
@@ 5.3  Symmetric 1:1 electrolyte: Transition to incompressible behavior with increasing Bulk modulus
The physical implications of this density variation are profound and directly impact the electrolyte’s ability to store charge. As discussed in Section 5.7, the enhanced ionic accommodation in compressible systems at high potentials, evident in Fig. 7, is the very mechanism that boosts the double-layer capacitance, $C_{dl}$. Conversely, as $K$ increases and the system becomes incompressible, the number density is constrained near $n^{0}$, reducing the space charge distribution and consequently lowering the overall $C_{dl}$.
@@ 5.4  Ionic distributions in annular battery designs at equilibrium
Figure 8: Equilibrium distribution of ions (cation and anion), potential, and total number density on full domain (left) and cross-section along $y=0$, $1\leq x\leq 2$ (right) in an annular battery design in Example 5.4, with an applied potential of 1.0-1.0 unit on the inner boundary and $1.0$ unit on the outer boundary. Parameter settings are provided in Table 3. Boundary conditions are described in Section 3.2.
@@ 5.7  $K$-dependent double layer capacitance curves
The results show that the total electrostatic charge stored in the system increases monotonically as $\varphi^{\mathrm{Left}}$ is varied from 0 to 80 for all values of the compressibility parameter $K$. As $K$ is increased, the total charge decreases and approaches the value corresponding to the incompressible mixture, in agreement with theoretical expectations. This behavior confirms that decreasing the compressibility of the system reduces the overall charge accumulation in the double layer. Also, near the electrodes, one ionic species becomes strongly depleted due to the applied potential, resulting in a negligible contribution of that species to the total charge. Consequently, the double layer behavior in this region is primarily governed by the counter-ions that accumulate close to the electrode surface.
The differential double layer capacitance, $C_{dl}$, exhibits a similar dependence on the compressibility parameter $K$, but shows a distinct behavior with respect to the applied potential. For high values of $K$ and the incompressible case, the capacitance exhibits the characteristic camel-shaped profile. For lower values of $K$, the capacitance initially increases with $\varphi^{\mathrm{Left}}$, reaches a shallow local maximum, then decreases slightly before increasing again at higher potentials. This non-monotonic behavior is associated with changes in the charge distribution and may be influenced by the additional density variations permitted in the compressible electrolyte (see Section 5.3). As $K$ increases, the overall magnitude of $C_{dl}$ decreases and approaches the incompressible limit, indicating that less compressible systems (high values of $K$) have a reduced ability to store charge within the double layer. These results highlight that electrolyte compressibility can play an important role in regimes where density variations significantly affect interfacial charge distributions.
@@ Acknowledgments
Ankur acknowledges the funding through the Advanced Research Opportunities Program (AROP), Aachen University, Germany. S.S. is funded by German Research Foundation (DFG) under Research Unit FOR5409: ”Structure-Preserving Numerical Methods for Bulk- and Interface-Coupling of Heterogeneous Models (SNuBIC)”. The authors also sincerely thank Prof. Manuel Torrilhon (Head of the ACoM Lab, RWTH Aachen University, Germany) for his valuable guidance and intellectual input throughout the entire AROP project carried out by Ankur at the ACoM Lab.